Please explain steps thoroughly of how one would solve this using technology. Is there also a way to avoid using the standard normal table or is it necessary? thanks!A particular fruits weights are normally distributed, with a mean of 456 grams and a standard deviation of 7 grams. If you p
Answer #1

Solution:-

Given that,

mean = \mu = 456

standard deviation = \sigma = 7

n = 21

\mu\bar x =\mu = 456

\sigma\bar x = \sigma / \sqrt n = 7/ \sqrt 21 = 1.5275

P(Z > z ) = 0.16

1- P(z < z) =0.16

P(z < z) = 1-0.16 = 0.84

z = 0.99

Using z-score formula,

\bar x = z * \sigma \bar x + \mu \bar x  = 0.99*1.5275 + 456 = 457.51

\bar x = 458 gram